# ALMA chemical survey of disk-outflow sources in Taurus (ALMA-DOT)

### v2

@@ 5 Discussion
As mentioned in the Introduction, different scenarios have been proposed for the interpretation of the recent observations of layered molecular winds. These scenarios are, in principle, also applicable to the HL Tau case. Ellipses in channel maps and transverse PVs, as well as an increase in velocity in longitudinal PVs are predicted by models of magnetized disk winds (Ray & Ferreira, 2021; de Valon et al., 2022), models of wide-angle winds driven swept-up shells (Lee et al., 2000; Zhang et al., 2019), unified models with a jet-bearing wide-angle X-wind (Shang et al., 2023; Ai et al., 2024). On the other hand, flow rotation is expected in both magnetically driven winds, which extract angular momentum from the disk, and thermal photoevaporative winds (Alexander et al., 2014), which conserve their initial angular momentum, but do not slow down the rotation of the disk. The layered wind structure could also be due to the interaction with a pulsed axial jet (e.g., Rabenanahary et al. (2022)). Each of these possibilities deserves a dedicated analysis. In the following, we focus on the scenario of extended MHD disk wind (DW), deferring the comparison of the observed trends with other classes of models to forthcoming papers.
5.1 Basics of the MHD Disk-Wind theory.
According to the Disk-Wind theory, a young star-disk system is embedded in a large-scale hourglass-shaped, twisted magnetic field inherited from the collapse of a magnetized rotating cloud. In this configuration, and provided sufficient ionization, a wind can be launched by the action of magnetic effects from the surface of the rotating disk. The outflow carries away vertically angular momentum, and the consequent braking of the disk rotation allows accretion through the disk to occur. Self-similar steady and axisymmetric DW solutions in ideal MHD conditions have proven to be successful in reproducing the observed properties of the axial atomic jets (see, e.g., Frank et al. 2014; Pascucci et al. 2023). The wind structure predicted by such models, that is, nested rotating gas shells, with higher poloidal velocities along the inner shells, is actually observed also in the outer CO outflows, as shown above for the HL Tau case. However, to properly describe global disk winds from an extended launch region, the models have to include non-ideal effects in the disk. Among them, ambipolar diffusion (AD) between neutrals and ions becomes dominant when ionization is scarce (e.g. Pinto et al. (2008)). Magnetized winds also arise in this case, as described, for instance, in the review by Lesur et al. (2023).
An important quantity of the DW theory is the parameter , which represents the ratio between the extracted and the initial specific angular momentum. Following the definition of Blandford & Payne (1982), can be expressed as:
@@
The importance of gas pressure versus magnetic effects in the disk can be parametrized by the plasma at the midplane, defined as , where is the sound speed and are the mean density and vertical magnetic field threading the disk. Unfortunately, the magnetic field amplitude in the disk is not available from observations. However, according to theoretical studies for self-similar disk wind solutions, is expected to increase with smaller (e.g. Jacquemin-Ide et al. 2019; Lesur 2021). It follows that the determination of from the observations is crucial not only to identify the properties of the wind, but also to constrain the importance of magnetic effects in the disk.
Self-similar DW solutions from fully turbulent disks have been found for a wide range of values at both high and low (Ferreira, 1997; Casse & Ferreira, 2000a; Jacquemin-Ide et al., 2019; Zimniak et al., 2024). On the other hand, in disk regions where the ionization is quenched in the midplane, as expected in the so-called ’dead zone’ of the disk, solutions are dominated by AD, and low values of are expected, up to about 3 (Lesur, 2021). Recent results show that the dead zone could actually start closer than 1 au from the star in T Tauri disks (Flock et al., 2019), and extend to the outer disk regions of interest here.
@@
5.2 Derivation of the shell footpoints in the Disk Wind scenario.
Regardless of the conditions adopted for the disk physics, the wind can be treated in ideal MHD. Assuming that the flow is a cold stationary and axisymmetric disk wind, from the measurements described in Sect. 4.1 one can infer the so-called footpoint radius , that is, the radius of the disk annulus from which the wind shell originates. Following the method introduced in Bacciotti et al. (2002) and refined in Anderson et al. (2003), can be found as the real solution of the following third-order equation:
where is the gravitation constant and the mass of the central source. The equation is valid if the enthalpy and the gravitational force exerted by the star are negligible with respect to the kinetic terms at the height above the disk where the wind parameters are measured. These conditions are fulfilled for cold CO molecular flows with poloidal velocities around 10 km sobserved at more than 500 au from a star of solar type, conditions met in our case. In addition, it is assumed that the flow propagates freely. Actually, the environment around HL Tau has a very complex structure, and before reaching the observed location the outflow may have interacted with several elements disturbing the steady propagation, such as accretion streamers (Garufi et al., 2022), winds inflated by nearby stars (Welch et al., 2000), surrounding cavities, and with the lateral wings of the bow-shocks of the inner wind components. However, the regular geometry observed on large scales supports the assumption that the kinematical information survives the occurring perturbations. We note that the observed emission in the CO 2-1 line is seen just because of the density contrast, even if the flow is perfectly laminar, as the emission does not require shock heating to be produced, since the energy of the upper level is only 17 K. However, the presence of external perturbations cannot be excluded a priori, and our simple model should be regarded as a first approximation to the real situation. The investigation of such effects has been deferred to forthcoming work.
@@
Interestingly, the footpoint radii of each of the B1a, B1b, and B2 shells cluster around the radius of one of three consecutive rings in the dust emission distribution in the disk, located at 58, 72 and 86 au from the source, respectively (labeled B4, B5 and B6 in ALMA Partnership et al. (2015))555Offsets from the source in ALMA Partnership et al. (2015) have been rescaled to a distance from HL Tau of 147.3 pc, following Galli et al. (2018).. However, note that the uncertainty of 20% in the determined footpoint radius (estimated by testing the sensitivity of the calculation in Eq. 3 to variations in the input parameters) also covers the positions of adjacent disk substructures.
@@
An equivalent derivation of the shell footpoints is illustrated by the diagnostic diagram of Fig. 12, constructed following the procedure described in Ferreira et al. (2006). Under the same assumptions, in fact, it can be shown that and for a given streamline depend only on the two parameters and . The location of the measured () values on the grid of theoretical curves drawn in the diagram gives estimates of the radius of the footpoint r0 and of the parameter . In HL Tau, for each individual shell the points lie along or near a curve at a defined value of , corresponding to the position of one of the disk rings B4, B5, and B6, and progress toward higher values of as the distance from the source increases. As discussed in the next section and in Appendix G, this trend is consistent with the one expected in a DW scenario.
@@
We warn that the positions of the wind footpoints should be compared with the distribution of disk substructures in the gas, rather than in the dust, and these may not coincide. It is well known that bumps in the gas pressure act as dust traps, but the average shift of the peak positions of gas pressure and dust emission depends on the model adopted (see, e.g. Riols et al. 2020; Hu et al. 2022). Yen et al. (2019a) reports profiles of gas emission in the HL Tau disk in the molecular lines of HCO+(3-2) and HCO+(1-0). The profiles are quite shallow, with no clear substructure detected, but it is interesting to note that the footpoint radii of the B-shells are in the vicinity of peaks in the HCO+(3-2)/HCO+(1-0) ratio (see their Fig. 9), which, under the assumption of constant temperature, can be considered proportional to the gas density. According to this comparison, therefore, the wind shells appear to arise from gas rings. However, the association of the footpoint locations with rings remains tentative at this stage, as a more reliable identification would require higher precision in the measurements of the gas tracers.
@@
The derived footpoint radii are larger than our spatial resolution limit ( 40 au). In fact, with reference to Fig. 14, the base of the red-shifted wind appears resolved in the channel maps at low velocities between VLSR= 7.4 and 9.4 km s. However, the emission pattern is not clear in these maps, as contributions at low velocity and close to the star may come both from the outflow base and from parts of the envelope, possibly from accreting streamers, or from some parts of the disk. Nevertheless, as discussed in Appendix A, this velocity interval is consistent with that in which, under the DW hypothesis, we expect to resolve the flow base in these observations. In the channel maps at VLSR km sthe wind base is not resolved on the SE side. At this velocity the flow appears to originate within a region inside the resolution limit of 40 au, it propagates initially with a large opening angle, up to a projected distance of about 0.”5 from the disk, and then it appears to recollimate. This trend is again consistent with the DW scenario, following which the inner shells coming from an unresolved footpoint are subject to a faster acceleration, have an initial large opening angle, which then reduces to the collimated value (cf. Fig. 13, first panel). We could not apply the same analysis to the NW side as the peak of the disk emission at these velocities covers the signal from the wind.
@@ 5.3 Variation of parameter in the wind shells
In Section 5.2, we determined at each distance sampled from the source through either its direct calculation as in Footnote 3, or graphically as in the diagram of Fig. 12. For all points relative to HL Tau is within the range = 2.5-4.0, and for each shell, the points relative to PV⟂ in subsequent offsets follow curves at constant , and are found at higher values of as the distance from the source increases. We note that the same behavior is shown by the points corresponding to theoretical DW models, in a zone in which the asymptotic regime has not yet been reached (see Appendix G). This appears to indicate that also in HL Tau we are probing a transition region in which the magnetic field is still transferring angular momentum to the matter. However, the increasing vicinity of the points for higher appears to indicate that the distances probed are not far from the asymptotic regime. This point should be farther away for the outer streamlines than for the inner ones, due to the self-similar nature of the flow (Ferreira et al., 2006). This is in line with the fact that a similar value of is attained closer to the star for the B1 shells than for B2 (sampled at a larger distance).
We also show in Fig. 12 the values of and derived for the shells identified in the HH30 CO outflow (Louvet et al., 2018; López-Vázquez et al., 2024) and in the bright conical DG Tau B CO outflow (de Valon et al., 2022). Interestingly, the and measurements in HH 30 are consistent with a curve at constant , while the points in DG Tau B follow a curve at constant . We caution, however, that the choice of the sampling points of the tomography in DG Tau B differs from ours (see Appendix E). In fact, the behavior determined in DG Tau B is expected in the case in which the tomography is probing a surface of varying in a homogeneous disk wind. Averaging our () values along each HL Tau shell would lead to a trend similar to that observed in DG Tau B, although with larger values of and .
5.4 Comparison with self-similar MHD disk-wind solutions.
@@
We do not attempt to explore the full range of possible self-similar solutions, but we show expectations from only three wind models that together encompass the range of values compatible with the observations. Turbulence-dominated disk models should be used for MHD winds anchored below a few au, and AD-dominated models further out. However, in both cases, the MHD wind is described by the same set of ideal MHD equations, and what differs is only the physical conditions at the jet launching point.
We use two solutions, L1 and L2, corresponding to cold disk winds arising from an AD dominated disk (Lesur, 2021), with, respectively, and 63 and corresponding to = 1.56 and = 2.76. Both solutions assume an ambipolar diffusion parameter = 1 and no Ohmic diffusion. The disk and its atmosphere were assumed to be cylindrically isothermal (sound speed ), with a disk aspect ratio = 0.1 (where is the geometric thickness of the disk , with the Keplerian angular velocity). The value achieved by solution L2 is the highest published so far for AD-dominated disks.666The solutions in Lesur (2021) used in our work are freely available at https://github.com/glesur/PPDwind.
In order to investigate the comparison with an MHD wind with a higher magnetic lever arm, we computed a warm solution corresponding to the case of a fully turbulent disk, similar to the one presented in Tabone et al. (2020) to reproduce the ALMA observations of the rotating SO wind in HH212. Our solution N5 describes a wind that arises from a thinner disk with = 0.01, 10, an effective magnetic Prandtl number of 1, a turbulence parameter =2.3, a turbulence anisotropy = 4.2 and no turbulent pressure (=0) (Zimniak et al., 2024). Despite its large magnetic field, the ejection efficiency is high and the resulting outflow is dense. It achieves a magnetic lever arm 5.64 thanks to additional heating at the base of the wind (see the description in Casse & Ferreira 2000b). This heating is localized at the disk surface, and requires only 2% (=0.02, see Eq. 25 in that work) of the dissipated power within the disk. Further up, the wind becomes adiabatic.
The comparison with the models is illustrated in Fig. 13. The points relative to the three B shells, with the same color code as in Fig. 10, describe the wind parameters normalized by their respective average footpoint radius and by the corresponding Keplerian velocity . The bottom curves are from the wind model from the AD-dominated disk region in Lesur (2021), with curve L1 (orange) corresponding to the solution for , =1.56, and curve L2 (blue) for , =2.76. The upper curve N5 (purple) corresponds to the model of wind from a fully turbulent disk region that provides =5.64 (cf. Zimniak et al. 2024).
@@
This may indicate that the plasma in the outer regions of the HL Tau disk is smaller (that is, the magnetic effects are more important) than commonly assumed for this zone in protoplanetary disks. A stronger field produces stronger vertical compression and reduced wind density, leading to a larger . However, other effects enter the wind acceleration mechanism, such as the aspect ratio of the disk, the existence of heating at the disk surface, the amount of effective turbulence, the degree of diffusion anisotropy, and the amount of ambipolar diffusion present in the disk. Regarding the last point, the outer region of the disk might possibly present a level of ambipolar diffusion consistently higher than that assumed in solutions of type L, related to an unexpectedly low level of ionization. This might be due to the screening of cosmic rays (the dominant ionization agent at 50 au) by magnetic mirroring from the wind, as proposed in Cleeves et al. (2015) for the TW Hya disk from chemical arguments. Reducing the irradiation of these particles could thus reduce the mass loading on the field lines and increase (Lesur, 2021). An AD-dominated scenario, leading to a laminar-type configuration in the region where rings/gaps are observed, would also be in line with the fact that the rings appear to be remarkably axisymmetric in HL Tau, pointing towards a low level of turbulence in the plane of the disk (Pinte et al., 2016).
@@
In summary, the physical properties of the molecular CO wind from HL Tau appear to be consistent with a description in terms of a magnetized wind originating from a large region of the disk, up to 90 au and possibly beyond. The idea that not only the axial atomic jet originated in the first au from the star, but also the outer wind may be magnetic in nature implies that feedback from flows onto the disk may occur on a large scale, which can have important consequences for models of disk evolution and planet formation. First of all, such winds may be able to sustain accretion along the disk surface at large distances from the star, in regions where the disk is internally weakly ionized and MHD turbulence, and hence viscous accretion, is suppressed. However, other effects may come into play. As discussed in Pascucci et al. (2025), for example, magnetically driven winds can slow the drift of solids in the direction of the star and affect the migration of planetary embryos.
5.5 Implications of the observed shell separation.

### v3

@@ 5 Discussion
As mentioned in the Introduction, different scenarios have been proposed for the interpretation of the recent observations of layered molecular winds. These scenarios are, in principle, also applicable to the HL Tau case. Ellipses in channel maps and transverse PVs, as well as an increase in velocity in longitudinal PVs are predicted by models of magnetized disk winds (Ray21; DeValon22), models of wide-angle winds driven swept-up shells (Lee00; Zhang19), unified models with a jet-bearing wide-angle X-wind (Shang23; Ai24). On the other hand, flow rotation is expected in both magnetically driven winds, which extract angular momentum from the disk, and thermal photoevaporative winds (Alexander14), which conserve their initial angular momentum, but do not slow down the rotation of the disk. The layered wind structure could also be due to the interaction with a pulsed axial jet (e.g., Rabenanahary22). Each of these possibilities deserves a dedicated analysis. In the following, we focus on the scenario of extended MHD disk wind (DW), deferring the comparison of the observed trends with other classes of models to forthcoming papers.
5.1 Basics of the MHD disk-wind theory
According to disk-wind theory, a young star-disk system is embedded in a large-scale hourglass-shaped, twisted magnetic field inherited from the collapse of a magnetized rotating cloud. In this configuration, and provided sufficient ionization, a wind can be launched by the action of magnetic effects from the surface of the rotating disk. The outflow carries away vertically angular momentum, and the consequent braking of the disk rotation allows accretion through the disk to occur. Self-similar steady and axisymmetric DW solutions in ideal MHD conditions have proven to be successful in reproducing the observed properties of the axial atomic jets (see, e.g., Frank14; Pascucci23). The wind structure predicted by such models, that is, nested rotating gas shells with higher poloidal velocities along the inner shells, is actually observed also in the outer CO outflows, as shown above for the HL Tau case. However, to properly describe global disk winds from an extended launch region, the models have to include non-ideal effects in the disk. Among them, ambipolar diffusion (AD) between neutrals and ions becomes dominant when ionization is scarce (e.g. Pinto08). Magnetized winds also arise in this case, as described, for instance, in the review by Lesur23.
An important quantity of the DW theory is the parameter , which represents the ratio between the extracted and the initial specific angular momentum. Following the definition of Blandford-Payne82, can be expressed as
@@ 5.1 Basics of the MHD disk-wind theory
The importance of gas pressure versus magnetic effects in the disk can be parametrized by the plasma at the midplane, defined as , where is the sound speed and are the mean density and vertical magnetic field threading the disk. Unfortunately, the magnetic field amplitude in the disk is not available from observations. However, according to theoretical studies for self-similar disk wind solutions, is expected to increase with smaller (e.g., Jacquemin-Ide19; Lesur21). It follows that the determination of from the observations is crucial not only to identify the properties of the wind, but also to constrain the importance of magnetic effects in the disk.
Self-similar DW solutions from fully turbulent disks have been found for a wide range of values at both high and low (Ferreira97; CasseFer00; Jacquemin-Ide19; Zimniak24). On the other hand, in disk regions where the ionization is quenched in the midplane, as expected in the so-called ’dead zone’ of the disk, solutions are dominated by AD, and low values of are expected, up to about 3 (Lesur21). Recent results show that the dead zone could actually start closer than 1 au from the star in T Tauri disks (Flock19), and extend to the outer disk regions of interest here.
@@
5.2 Derivation of the shell footpoints in the disk wind scenario
Regardless of the conditions adopted for the disk physics, the wind can be treated in ideal MHD. Assuming that the flow is a cold stationary and axisymmetric disk wind, from the measurements described in Sect. 4.1 one can infer the so-called footpoint radius , that is, the radius of the disk annulus from which the wind shell originates. Following the method introduced in Bacciotti02 and refined in Anderson03, can be found as the real solution of the following third-order equation:
Here is the gravitation constant and the mass of the central source. The equation is valid if the enthalpy and the gravitational force exerted by the star are negligible with respect to the kinetic terms at the height above the disk where the wind parameters are measured. These conditions are fulfilled for cold CO molecular flows with poloidal velocities around 10 km sobserved at more than 500 au from a star of solar type, conditions met in our case. In addition, it is assumed that the flow propagates freely. Actually, the environment around HL Tau has a very complex structure, and before reaching the observed location the outflow may have interacted with several elements disturbing the steady propagation, such as accretion streamers (Garufi22), winds inflated by nearby stars (Welch00), surrounding cavities, and with the lateral wings of the bow-shocks of the inner wind components. However, the regular geometry observed on large scales supports the assumption that the kinematical information survives the occurring perturbations. We note that the observed emission in the CO 2-1 line is seen just because of the density contrast, even if the flow is perfectly laminar, as the emission does not require shock heating to be produced, since the energy of the upper level is only 17 K. However, the presence of external perturbations cannot be excluded a priori, and our simple model should be regarded as a first approximation to the real situation. The investigation of such effects has been deferred to forthcoming work.
@@ 5.2 Derivation of the shell footpoints in the disk wind scenario
Interestingly, the footpoint radii of the B1a, B1b, and B2 shells are found to correspond to the radii of three consecutive rings in the dust emission distribution in the disk, located at 58, 72, and 86 au from the source, respectively (labeled B4, B5, and B6 in ALMA15).555The position of the rings on the disk surface given in ALMA15 were rescaled for a distance from HL Tau of 147.3 pc, following Galli18. However, we note that the uncertainty of 20% in the determined footpoint radius (estimated by testing the sensitivity of the calculation in Eq. 3 to variations in the input parameters) also covers the positions of adjacent disk substructures.
@@
An equivalent derivation of the shell footpoints is illustrated by the diagnostic diagram of Fig. 12, constructed following the procedure described in Ferreira06. Under the same assumptions, in fact, it can be shown that and for a given streamline depend only on the two parameters and . The location of the measured () values on the grid of theoretical curves drawn in the diagram gives estimates of the radius of the footpoint r0 and of the parameter . In HL Tau, for each individual shell the points lie along or near a curve at a defined value of , corresponding to the position of one of the disk rings B4, B5, and B6, and progress toward higher values of as the distance from the source increases. As discussed in the next section and in Appendix G, this trend is consistent with the one expected in a DW scenario.
@@
We warn that the positions of the wind footpoints should be compared with the distribution of disk substructures in the gas, rather than in the dust, and these may not coincide. It is well known that bumps in gas pressure act as dust traps, but the average shift of the peak positions of gas pressure and dust emission depends on the model adopted (see, e.g. Riols20; Hu22). Yen19a reports profiles of gas emission in the HL Tau disk in the molecular lines of HCO+(3-2) and HCO+(1-0). The profiles are quite shallow, with no clear substructure detected, but it is interesting to note that the footpoint radii of the B-shells are in the vicinity of peaks in the HCO+(3-2)/HCO+(1-0) ratio (see their Fig. 9), which, under the assumption of constant temperature, can be considered proportional to the gas density. According to this comparison, therefore, the wind shells appear to arise from gas rings. However, the association of the footpoint locations with rings in the gas distribution remains tentative at this stage, as a more reliable identification would require higher precision in the measurements of the gas tracers.
@@
The derived footpoint radii are larger than our spatial resolution limit, which is about 40 au. With reference to Fig. 14, the base of the redshifted wind appears resolved in the channel maps at low velocities between VLSR= 7.4 and 9.4 km s. However, the emission pattern is not clear in these maps, as contributions at low velocity and close to the star may come both from the outflow base and from parts of the envelope, possibly from accreting streamers, or from some parts of the disk. Nevertheless, as discussed in Appendix A, this velocity interval is consistent with that in which, under the DW hypothesis, we expect to resolve the flow base in these observations. In the channel maps at VLSR km sthe wind base is not resolved on the SE side. At this velocity the flow appears to originate within a region inside the resolution limit of 40 au, it propagates initially with a large opening angle, up to a projected distance of about 0.”5 from the disk, and then it appears to recollimate. This trend is again consistent with the DW scenario, following which the inner shells coming from an unresolved footpoint are subject to a faster acceleration, have an initial large opening angle, which then reduces to the collimated value (cf. Fig. 13, first panel). We could not apply the same analysis to the NW side as the peak of the disk emission at these velocities covers the signal from the wind.
@@ 5.3 Variation of parameter in the wind shells
In Section 5.2, we determined at each sampled distance from the source either through its direct calculation, as in Footnote 3, or graphically as in the diagram of Fig. 12. For all points relative to HL Tau is within the range = 2.5 - 4.0, and for each shell, the points relative to PV⟂ in subsequent offsets follow curves at constant , and are found at higher values of as the distance from the source increases. We note that the same behavior is shown by the points corresponding to theoretical DW models, in a zone in which the asymptotic regime has not yet been reached (see Appendix G). This appears to indicate that also in HL Tau we are probing a transition region in which the magnetic field is still transferring angular momentum to the matter. However, the increasing vicinity of the points for higher suggests that the distances probed are not far from the asymptotic regime. This point should be farther away for the outer streamlines than for the inner ones, due to the self-similar nature of the flow (Ferreira06). This is in line with the fact that a similar value of is attained closer to the star for the B1 shells than for B2 (sampled at a larger distance).
We also show in Fig. 12 the values of and derived for the shells identified in the CO outflow of HH30 (Louvet18; Lopez24) and in the bright conical CO outflow of DG Tau B (DeValon22). Interestingly, the and measurements in HH 30 are consistent with a curve at constant , while the points in DG Tau B follow a curve at constant . We caution, however, that the choice of the sampling points of the tomography in DG Tau B differs from ours (see Appendix E). In fact, the behavior determined in DG Tau B is expected in the case in which the tomography is probing a surface of varying in a homogeneous disk wind. Averaging our () values along each HL Tau shell would lead to a trend similar to that observed in DG Tau B, although with higher values of and .
5.4 Comparison with self-similar MHD disk-wind solutions
@@
We did not attempt to explore the full range of possible self-similar solutions, but we show expectations from only three wind models that together encompass the range of values compatible with the observations. Turbulence-dominated disk models should be used for MHD winds anchored below a few au, and AD-dominated models farther out. However, in both cases, the MHD wind is described by the same set of ideal MHD equations, and what differs is only the physical conditions at the jet launching point.
We used two solutions, L1 and L2, corresponding to cold disk winds arising from an AD dominated disk (Lesur21), with, respectively, and 63 and corresponding to = 1.56 and = 2.76. Both solutions assume an ambipolar diffusion parameter = 1 and no Ohmic diffusion. The disk and its atmosphere were assumed to be cylindrically isothermal (sound speed ), with a disk aspect ratio = 0.1, where is the geometric thickness of the disk (, with the Keplerian angular velocity). The value achieved by solution L2 is the highest published so far for AD-dominated disks.666The solutions in Lesur21 used in our work are freely available at https://github.com/glesur/PPDwind.
In order to investigate the comparison with a MHD wind with a higher magnetic lever arm, we computed a warm solution corresponding to the case of a fully turbulent disk, similar to the one presented in Tabone20 to reproduce the ALMA observations of the rotating SO wind in HH212. Our solution N5 describes a wind that arises from a thinner disk with = 0.01, 10, an effective magnetic Prandtl number of 1, a turbulence parameter =2.3, a turbulence anisotropy = 4.2, and no turbulent pressure (=0) (Zimniak24). Despite its large magnetic field, the ejection efficiency is high and the resulting outflow is dense. It achieves a magnetic lever arm 5.64 thanks to additional heating at the base of the wind (see the description in CasseFer00b). This heating is localized at the disk surface and requires only 2% (=0.02, see Eq. 25 in that work) of the dissipated power within the disk. Farther up, the wind becomes adiabatic.
The comparison with the models is illustrated in Fig. 13. The points relative to the three B shells, with the same color code as in Fig. 10, describe the wind parameters normalized by their respective average footpoint radius and by the corresponding Keplerian velocity . The bottom curves are from the wind model from the AD-dominated disk region in Lesur21, with curve L1 (orange) corresponding to the solution for , =1.56, and curve L2 (blue) for , =2.76. The upper curve N5 (purple) corresponds to the model of wind from a fully turbulent disk region that provides =5.64 (cf. Zimniak24).
@@
This may indicate that the plasma in the outer regions of the HL Tau disk is smaller (that is, the magnetic effects are more important) than commonly assumed for this zone in protoplanetary disks. A stronger field produces stronger vertical compression and reduced wind density, leading to a larger . However, other effects enter the wind acceleration mechanism, such as the aspect ratio of the disk, the existence of heating at the disk surface, the amount of effective turbulence, the degree of diffusion anisotropy, and the amount of ambipolar diffusion present in the disk. Regarding the last point, the outer region of the disk might possibly present a level of ambipolar diffusion consistently higher than that assumed in solutions of type L, related to an unexpectedly low level of ionization. This might be due to the screening of cosmic rays (the dominant ionization agent at 50 au) by magnetic mirroring from the wind, as proposed in Cleeves15 for the TW Hya disk from chemical arguments. Reducing the irradiation of these particles could thus reduce the mass loading on the field lines and increase (Lesur21). An AD-dominated scenario, leading to a laminar-type configuration in the region where rings/gaps are observed, would also be in line with the fact that the rings appear to be remarkably axisymmetric in HL Tau, pointing toward a low level of turbulence in the plane of the disk (Pinte16).
@@
In summary, the physical properties of the molecular CO wind from HL Tau appear to be consistent with a description in terms of a magnetized wind originating from a large region of the disk, up to 90 au and possibly beyond. The idea that not only the axial atomic jet originated in the first au from the star, but also the outer wind may be magnetic in nature implies that feedback from flows onto the disk may occur on a large scale. This may have important consequences for models of disk evolution and planet formation. First of all, such winds may be able to sustain accretion along the disk surface at large distances from the star, in regions where the disk is internally weakly ionized and MHD turbulence, and hence viscous accretion, is suppressed. Other effects may also come into play. As discussed in Pascucci24, for example, magnetically driven winds can slow the drift of solids in the direction of the star and affect the migration of planetary embryos.
5.5 Implications of the observed shell separation